# If Probable, Then Acceptable? Understanding Conditional Acceptability Judgments in Large Language Models

### v1

@@ 4 Results
Despite these differences, the overall mean ratings by relation type are similar. Human means for the irrelevant and negative conditions (IRR: 0.250 519 1-0.250\,519\,1, NEG: 0.262 527 3-0.262\,527\,3) closely match the ones of all LLMs, which range from 0.37-0.37 to 0.20-0.20 for IRR and 0.27-0.27 to 0.19-0.19 for NEG (see Table 8 in the Appendix). The biggest difference occurs in the positive condition (humans POS 0.192 213 10.192\,213\,1), where most LLMs fall short. Qwen 72B is the only LLM that somewhat approximates this high mean across prompting techniques (vanilla: 0.160 312 50.160\,312\,5, few-shot: 0.152 145 80.152\,145\,8, CoT: 0.139 395 80.139\,395\,8). In terms of the influence of predictors, both humans and LLMs show strong effects of conditional probability and relation type. For humans, ANOVA results (see Table 19 in the Appendix) indicate significant main effects of probability and relevance, as well as their interaction ($p<0.0001$), but no effect of judgment metric. This stands in contrast to LLMs, where the smaller models and all few-shot and CoT models consistently show modest but significant effects of the judgment type. In terms of slope comparisons, human pairwise interactions across relation types are all highly significant ($p<0.0001$), indicating robust modulation by relevance. This pattern most closely resembles Qwen 7B (vanilla), Llama 70B (few-shot), and Qwen 7B (few-shot), although the LLMs generally produce weaker effects and differ in direction: humans show steeper slopes for positively related conditionals, whereas LLMs sometimes favor negative ones. Finally, estimated marginal means comparisons show that humans distinguish between relation types at all levels of conditional probability; except between irrelevant and negative conditionals at low probability ($p<0.0001$ for all other contrasts, see Table 20). This pattern is mirrored in the Qwen models and partially in the Llama models, which also exhibit collapsed distinctions at low probability levels.
@@ 6 Conclusion
In this paper, we investigate how large language models assess conditional statements, a task that has long been debated in the context of human reasoning, but remains underexplored for LLMs. Specifically, we examine whether LLMs rely on conditional probability, relation type, or a combination of both when judging the acceptability of conditionals. Our findings suggest that both factors contribute to LLM judgments, although the degree of this influence varies substantially across models and prompting strategies. Notably, while LLMs and humans perform generally similar, key distinctions emerge, suggesting that LLMs may not consistently integrate both semantic relation and probabilistic strength in the systematic way that humans do.
@@ 7 Limitations
As noted in Section 4, the few-shot and CoT prompting conditions appear to bias model sensitivity towards negatively related conditionals. Since this effect occurred in all few-shot and CoT variants, the random selection of prompt examples (see Appendix A.3) may have inadvertently biased model behavior. Given the well-documented prompt sensitivity of LLMs (Errica et al., 2024; Loya et al., 2023; Sclar et al., 2023), future work should take greater care in constructing prompt sets. One possible improvement could be to include a balanced set of few-shot examples, e.g., one for each combination of relation type and conditional probability level (resulting in nine examples total). This might help ensure that models are not inadvertently primed to favor particular judgment patterns.
@@ A.1 Prior Human Study
Relation / Prob. High–High High–Low Low–High Low–Low
Positive (POS) POS $\times$ High–High POS $\times$ High–Low POS $\times$ Low–High POS $\times$ Low–Low
Negative (NEG) NEG $\times$ High–High NEG $\times$ High–Low NEG $\times$ Low–High NEG $\times$ Low–Low
Irrelevant (IRR) IRR $\times$ High–High IRR $\times$ High–Low IRR $\times$ Low–High IRR $\times$ Low–Low
Table 3: Design matrix showing all 12 conditional statement configurations for a single scenario, based on 3 relation types (positive, negative, irrelevant) and 4 prior probability levels (high-high, high-low, low-high, low-low).
@@ A.2 Models and Prompts
This study employs four publicly available large language models of comparable size (787-8 billion parameters and 707270-72): Llama 3.1 (8B and 70B) and Qwen 2.5 (7B and 72B). All models are instruction-tuned variants designed to follow user prompts and generate coherent natural language responses.
@@
Llama-3.1-8B-Instruct 8.03B 128K
Llama-3.1-70B-Instruct 70.6B 128K
Qwen2.5-7B-Instruct 7.61B 128K
Qwen2.5-72B-Instruct 72.7B 128K
@@
Sentence probability Perplexity
Model Pearson Spearman Pearson Spearman
Llama 8B cond. prob. 0.15 0.25 -0.28 -0.32
Llama 8B if-prob. 0.19 0.25 -0.32 -0.32
Llama 70B cond. prob. 0.16 0.26 -0.28 -0.38
Llama 70B if-prob. 0.11 0.31 -0.23 -0.31
Qwen 7B cond. prob. 0.08 0.22 -0.17 -0.27
Qwen 7B if-prob. 0.07 0.20 -0.08 -0.27
Qwen 72B cond. prob. 0.16 0.33 -0.21 -0.36
Qwen 72B if-prob. 0.14 0.19 -0.19 -0.31
Table 6: Pearson and Spearman correlation of sentence probability and perplexity with probability judgments obtained through direct elicitation.
@@ A.3 Prompt Templates
Figure 8 shows the vanilla prompt used to elicit conditional probability estimates. Similar to the original study (Skovgaard-Olsen et al., 2016), the model is given a context, a suppositional assumption (e.g., “Suppose Nicole sunbathes on the beach.”), and a target sentence (e.g., “She will get sunburned.”). It is then asked how probable the target sentence is, given the assumption.
@@
A.4 Probability Elicitation Methods
As described in Section 3.1 and A.2, we directly prompt models to judge the (conditional) probability or acceptability of a conditional statement. While this method is well-established and has been used by existing studies (Tian et al., 2023; Yang et al., 2025), it is important to note that other methods exist for obtaining a model’s probability estimate for a given statement. For instance, it is possible to compute the sentence probability a model assigns to a given statement by leveraging its output probability distribution over the model’s vocabulary. While both direct elicitation and sentence probability have their respective advantages and disadvantages (as illustrated in Table 5), there are key differences between them that should be considered. Sentence probability quantifies how likely a model considers a given string. Notably, this captures surface form rather than semantics and does not necessarily reflect how likely the model deems the given statement to be true. In addition, pure sentence probability is highly sensitive to paraphrasing and tokenization and exhibits a notable length bias. In contrast, direct elicitation allows the model to be asked directly for the probability that a given statement is true. For instruct/chat models in particular, direct elicitation has been shown to be better calibrated than token likelihoods (Tian et al., 2023), as it leverages the model’s internal latent knowledge and reasoning.
To compare alternative methods, such as sentence probability, with our direct elicitation approach, we compute the sentence probability of both the conditional probability $P(B\mid A)$ and the if-probability P(If A, then B)P($\text${If A, then B}) (as defined in Section 3.1). In particular, for conditional probability, we use the following prompt $M$:
Context: <context> Given: <A> Consequence:
[…]
Context: <context> Statement:
Because sentence probability is highly dependent on sentence length, we additionally compute the model’s perplexity over the statements. To study how well these probability-based judgments correlate with those obtained via direct elicitation, we compute both Pearson and Spearman correlations for each metric. As shown in Table 6, we observe a weak positive correlation for sentence probability and a weak-to-moderate negative correlation for perplexity scores, underlining that these metrics are related but capture different aspects.
Model (Prompt) Conditional Probability If-Probability Acceptability
Llama-3.1-8B (Vanilla) 0.7685 0.7724 0.7529
Llama-3.1-70B (Vanilla) 0.9468 0.9344 0.9201
Qwen-2.5-7B (Vanilla) 0.9768 0.9716 0.9450
Qwen-2.5-72B (Vanilla) 0.9693 0.9673 0.9784
Llama-3.1-8B (Few-shot) 0.7930 0.7538 0.7596
Llama-3.1-70B (Few-shot) 0.9438 0.9556 0.8820
Qwen-2.5-7B (Few-shot) 0.9542 0.9556 0.9596
Qwen-2.5-72B (Few-shot) 0.9559 0.9437 0.9768
Llama-3.1-70B (CoT) 0.9014 0.8497 0.8757
Qwen-2.5-72B (CoT) 0.9331 0.9127 0.9661
Table 7: Intraclass correlation scores across five model runs for each score type, model and prompt type.
A.5 Consistency of Model Responses
Because LLM inference is non-deterministic, ensuring the consistency of the models’ numerical probability ratings is an important consideration. To address potential instabilities in model ratings, we perform repeated trials by prompting each model five times per data sample, as described in Section 3.2. We further evaluate the consistency of the models’ probability judgments across runs by computing the intraclass correlation of the five ratings per sample for each score type and model. As shown in Table 7, all models demonstrate good-to-excellent reliability across all score types and prompting strategies.
A.6 Linear Mixed-Effects Model
@@
Model Specification and Adaptation
Each linear mixed-effects model uses judgment, i.e. the judgments of conditional statements, as the dependent variable. A single model is fit per respondent (LLM or human) with the same structure, enabling later comparisons. Following best practices (Bates et al., 2015; Brown, 2021), we balance model complexity and convergence. Due to singularity issues with full random-slopes structures, we simplify the models by including random intercepts for both instance ID and scenario in the LLM linear mixed-effects models. The human mixed-effects model, however, is fitted with instance ID as a random slope instead, as for humans, this represents the individual participants and is expected to introduce much greater variation. This choice captures key variance across participants and model samples while retaining model stability.
A.7 ANOVA Tables
Tables 9 to 18 illustrate the results from the ANOVA analyses.

### v2

@@ 4 Results
Despite these differences, the overall mean ratings by relation type are similar. Human means for the irrelevant and negative conditions (IRR: 0.250 519 1-0.250\,519\,1, NEG: 0.262 527 3-0.262\,527\,3) closely match the ones of all LLMs, which range from 0.37-0.37 to 0.20-0.20 for IRR and 0.27-0.27 to 0.19-0.19 for NEG (see Table 8 in the Appendix). The biggest difference occurs in the positive condition (humans: 0.192 213 10.192\,213\,1), where most LLMs fall short. Both Llama models and the smaller Qwen variant barely exceed the midpoint (Llama 8B: $0.01$, 70B: $0.04$; Qwen: $0.06$). Qwen 72B is the only LLM that approximates this high mean across prompting techniques (vanilla: 0.160 312 50.160\,312\,5, few-shot: 0.152 145 80.152\,145\,8, CoT: 0.139 395 80.139\,395\,8). In terms of the influence of predictors, both humans and LLMs show strong effects of conditional probability and relation type. For humans, ANOVA results (see Table 19 in the Appendix) indicate significant main effects of probability and relevance, as well as their interaction ($p<0.0001$), but no effect of judgment metric. This stands in contrast to LLMs, where the smaller models and all few-shot and CoT models consistently show modest but significant effects of the judgment type. In terms of slope comparisons, human pairwise interactions across relation types are highly significant ($p<0.0001$), indicating robust modulation by relevance. This pattern most closely resembles Qwen 7B (vanilla), Llama 70B (few-shot), and Qwen 7B (few-shot), although the LLMs generally produce weaker effects and differ in direction: humans show steeper slopes for positively related conditionals, whereas LLMs sometimes favor negative ones. Finally, estimated marginal means comparisons show that humans distinguish between relation types at all levels of conditional probability; except between irrelevant and negative conditionals at low probability ($p<0.0001$ for all other contrasts, see Table 20). This pattern is mirrored in the Qwen models and partially in the Llama models, which also exhibit collapsed distinctions at low probability levels.
@@
Sentence Prob. Perplexity
$\mathbf{r}$ $\mathbf{\rho}$ $\mathbf{r}$ $\mathbf{\rho}$
Cond. Prob.
Llama 8B 0.15 0.25 -0.28 -0.32
Llama 70B 0.16 0.26 -0.28 -0.38
Qwen 7B 0.08 0.22 -0.17 -0.27
Qwen 72B 0.16 0.33 -0.21 -0.36
If-Prob.
Llama 8B 0.19 0.25 -0.32 -0.32
Llama 70B 0.11 0.31 -0.23 -0.31
Qwen 7B 0.07 0.20 -0.08 -0.27
Qwen 72B 0.14 0.19 -0.19 -0.31
Table 4: Pearson $\left(\mathbf{r}\right)$ and Spearman correlation $\left(\mathbf{\rho}\right)$ of sentence probability and perplexity with probability judgments obtained through direct elicitation.
4.2 Eliciting Probability Judgments
In our study, we directly prompt models to judge the probability or acceptability of a conditional statement. While this method is well-established and has been used in existing studies (Tian et al., 2023; Yang et al., 2025), there also exist other methods for eliciting ratings from LLMs. [^6] A discussion of direct elicitation and alternative methods is provided in Appendix A.4. For instance, we can compute a model’s sentence probability or its perplexity as an estimate of the probability it assigns to a given statement.
To compare sentence probabilities with our direct elicitation approach, we compute the model-assigned probabilities of both the conditional probability $P(B\mid A)$ and the if-probability P(“If A, then B”)P($\text${``If A, then B''}). In particular, for conditional probability, we use the following prompt $M$:
Context: <context> Given: <A> Consequence:
[…]
Context: <context> Statement:
Because sentence probability strongly depends on sentence length, we additionally compute the model’s perplexity over the statements. To study how well these probability-based judgments correlate with those obtained via direct elicitation, we compute both Pearson and Spearman correlations for each metric. As shown in Table 4, we observe a weak positive correlation for sentence probability and a weak-to-moderate negative correlation for perplexity scores, underlining that these metrics are related but ultimately capture different aspects.
@@ 6 Conclusion
In this paper, we investigate how LLMs assess conditional statements, a task that has long been debated in the context of human reasoning, but remains underexplored for LLMs. Specifically, we examine whether LLMs rely on conditional probability, relation type, or a combination of both when judging the acceptability of conditionals. Our findings suggest that both factors contribute to LLM judgments, although the degree of this influence varies substantially across models and prompting strategies. Notably, while LLMs and humans perform generally similar, key distinctions emerge, suggesting that LLMs may not consistently integrate both semantic relation and probabilistic strength in the systematic way that humans do.
@@ 7 Limitations
As noted in Section 4, the few-shot and CoT prompting conditions appear to bias model sensitivity towards negatively related conditionals. Since this effect occurred in all few-shot and CoT variants, the random selection of prompt examples (see Appendix A.3) may have inadvertently biased model behavior. Given the well-documented prompt sensitivity of LLMs (Errica et al., 2024; Loya et al., 2023; Sclar et al., 2023), future work should take greater care in constructing prompt sets. One possible improvement could be to include a more balanced set of few-shot examples, e.g., one for each combination of relation type and conditional probability level (resulting in nine examples total). This might help ensure that models are not inadvertently primed to favor particular judgment patterns.
@@ A.1 Prior Human Study
In total, this then results in 12 statements per scenario, each representing a unique combination of prior probability and relation type.
@@ A.2 Models and Prompts
This study employs four publicly available large language models of comparable size (787-8 and 707270-72 billion parameters): Llama 3.1 (8B and 70B) and Qwen 2.5 (7B and 72B). All models are instruction-tuned variants designed to follow user prompts and generate coherent natural language responses.
@@
Llama-3.1-8B-it 8.03B 128K
Llama-3.1-70B-it 70.6B 128K
Qwen2.5-7B-it 7.61B 128K
Qwen2.5-72B-it 72.7B 128K
@@ A.3 Prompt Templates
Figure 8 shows the vanilla prompt used to elicit conditional probability estimates. Similar to the original human study (Skovgaard-Olsen et al., 2016), the model is given a context, a suppositional assumption (e.g., “Suppose Nicole sunbathes on the beach.”), and a target sentence (e.g., “She will get sunburned.”). It is then asked how probable the target sentence is, given the assumption.
@@
Method How It Works Advantages Disadvantages
A.4 Discussion of Elicitation Methods
While we use direct elicitation to retrieve (conditional) probability and acceptability ratings (cf. Sections 3.1 and A.2), a well-established approach (e.g., Tian et al., 2023; Yang et al., 2025), other methods exist for obtaining a model’s probability estimate for a given statement. As an alternative to direct elicitation, we also compute the sentence probability a model assigns to a given statement by leveraging its output probability distribution over the model’s vocabulary (see Section 4.2).
While both direct elicitation and sentence probability have their respective advantages and disadvantages (as illustrated in Table 7), they bear key differences that should be considered. Sentence probability quantifies how probable a model considers a given string. This captures surface form rather than semantics and does not necessarily reflect how likely the model deems the statement to be true. Additionally, pure sentence probability is highly sensitive to paraphrasing and tokenization and exhibits a notable length bias. In contrast, direct elicitation allows the model to be asked directly for the probability that a given statement is true. For instruct/chat models in particular, direct elicitation has been shown to be better calibrated than token likelihoods (Tian et al., 2023), as it leverages the model’s internal latent knowledge and reasoning.
A.5 Linear Mixed-Effects Model
@@
Model Specification and Adaptation
Each linear mixed-effects model uses judgment, i.e., the judgments of conditional statements, as the dependent variable. A single model is fit per respondent (LLM or human) with the same structure, enabling later comparisons. Following best practices (Bates et al., 2015; Brown, 2021), we balance model complexity and convergence. Due to singularity issues with full random-slopes structures, we simplify the models by including random intercepts for both respondent (sampling iteration) and scenario in the LLM linear mixed-effects models. The human mixed-effects model, however, is fit with respondent as a random slope instead, as for humans, this represents the individual participants and is expected to introduce much greater variation. This choice captures key variance across participants and model samples while retaining model stability.
A.6 ANOVA Tables
Tables 9 to 18 illustrate the results from the ANOVA analyses.